OpenStudy (anonymous):

Exponents and logarithms and radical numbers...sooo simplify the expression nr.3

OpenStudy (anonymous):

\[\frac{ 3\sqrt{50} }{ 5\sqrt{27} }\]

OpenStudy (jdoe0001):

can you factor out 50? and 27? or so-called, simplify them?

OpenStudy (anonymous):

yes? or can you please show me how to do the whole thing @jdoe0001

OpenStudy (jdoe0001):

lemme simplify 50 and 27 so you see it  \(\bf 50 = 2 \times 5 \times 5 \implies 2 \times 5^2\\\quad \\ 27 = 3\times 3 \times 3 \implies 3^2 \times 3\\\\ \quad \\ \quad \\ \cfrac{ 3\sqrt{ 50 } }{ 5\sqrt{ 27 } } \implies \cfrac{ 3\sqrt{ 2 \times 5^2 } }{ 5\sqrt{ 3^2 \times 3 } }\)